Liquid chlorotribromomethane has a density of 2.71 g/mL. A barometer is constructed using chlorotribromomethane instead of mercury. If the atmospheric pressure is 0.950 atm, what is the height of the chlorotribromomethane column in the barometer in cm?

The density of liquid mercury is 13.6 g/mL. Answer is in cm
1
Expert's answer
2017-12-07T03:03:07-0500
For two liquids, A and B, that exert the same pressure, there is next expression: P=DA*hA=DB*hB
Since the conversion factor between mmHg and atmospheres is known, first express the pressure of 0.950 atmospheres in mmHg: 0.950 atm x (760 mmHg /atm) = 722 mmHg
Substituting A for chlorotribromomethane and В for mercury gives
hA = (DHg/DA)*hHg = (13.6g/mL / 2.71g/mL)x 722 mm = 3623 mm = 362.3 cm

Answer: 362.3 cm